Problem 2 Solve the following initial value problem (IVP). (1) =V2 xyy = y2 + x*e*? set u=2
Answer #1

Given that xyy= y² + x + x² ; y (1) = 1/2 => y = y² + xe = y= 4 + . et? put; wa > y = ux y= utxu from Equation ; we get 3